OpenStudy (gravierr):

I Really need help with 5 questions can someone please help me and can you show me your work so i could have a better understanding Thank You:) 1.Juan can clean a typical pool in 5 hours. Carl can clean a typical pool in 4 hours. How long should it take Juan and Carl working together to clean a typical pool? Show all your work. Leave your final answer as an integer or reduced fraction. 2. Simplify 5 * sqrt(8) / sqrt(40) by rationalizing the denominator 3. Simplify (3sqrt(5) + 2sqrt(6))^2. 4. Simplify the radical expression. Show all your steps. sqrt(192) – 4sqrt(27) I'd really appreci

OpenStudy (gravierr):

@ShadowLegendX

OpenStudy (shadowlegendx):

time = 5 hrs/typical pool ; rate = 1/5 , typical pool/hr time = 4 hrs/typical pool ; rate = 1/4 , typical pool/hr

OpenStudy (gravierr):

could you walk me through the problem Please

OpenStudy (shadowlegendx):

time = x hrs/typical pool ; 1/x , typical pool/hr

OpenStudy (shadowlegendx):

Our equation is rate + rate = together rate

OpenStudy (gravierr):

So i just add the 2 rates and thats my answer?

OpenStudy (shadowlegendx):

1/5 + 1/4 = 1/x Multiply whole equation by x 1/5x + 1/4x = 1 Multiply whole equation by 5 x + 5/4x = 5 Multiply whole equation by 4 4x + 5x = 20 9x = 20 x = 20/9 or 2.22222222222 = x x = how many hours it would take

OpenStudy (shadowlegendx):

Proof the above is correct, http://prntscr.com/djzktb

OpenStudy (gravierr):

Thank you very much

OpenStudy (shadowlegendx):

2. \[ 5 * \sqrt(8) / \sqrt(40) = 5 \times \frac{ \sqrt 8 }{ \sqrt 40 }\] \[5 \times \frac{ \sqrt 8 }{ \sqrt 40 } \rightarrow 5 \times \frac{ \sqrt 8 }{ \sqrt 40 } \times \frac{ \sqrt 40 }{ \sqrt 40 }\]

OpenStudy (gravierr):

I've got to go somewhere in 10 minutes so i will be back for sure and will medal and fan. Thank you for your help. There are only 4 more questions i would like help with

OpenStudy (shadowlegendx):

\[5 \times \frac{ \sqrt 320 }{ 40 }\]

OpenStudy (shadowlegendx):

We know that 8 x 40 = 320, we can start there with breaking down the sqrt of 320  \[5 \times \frac{ 8 \sqrt 5 }{ 40 }\]